### v4

# The Streaming Batch Model for Efficient and Fault-Tolerant Heterogeneous Execution

Abstract
While ML model training and inference are both GPU-intensive, CPU-based data processing is often the bottleneck. Distributed data processing systems based on the batch or stream processing models assume homogeneous resource requirements. They excel at CPU-based computation but either under-utilize heterogeneous resources or impose high overheads on failure and reconfiguration. We introduce the streaming batch model, a hybrid of the two models that enables efficient and fault-tolerant heterogeneous execution. The key idea is to execute one partition at a time to allow lineage-based recovery with dynamic resource allocation. This enables memory-efficient pipelining across heterogeneous resources, similar to stream processing, but also offers the elasticity and fault tolerance properties of batch processing. We present Ray Data, an implementation of the streaming batch model that improves throughput on heterogeneous batch inference pipelines by 3–8$\times$ compared to traditional batch and stream processing systems. When training Stable Diffusion, Ray Data matches the throughput of single-node ML data loaders while additionally leveraging distributed heterogeneous clusters to further improve training throughput by 31%.
11footnotetext: Work done while at Anyscale.
1 Introduction
Data processing is critical to machine learning applications. While model training and inference are both GPU-intensive, they also require significant I/O and CPU to load and preprocess datasets. CPU-based preprocessing is often the bottleneck in both training [30] and batch inference [22]. Meanwhile, as ML models evolve, the data processing functionality required has also become more diverse, spanning many modalities, transformations, and resource requirements [19, 41, 32].
There are numerous frameworks designed to scale out CPU-based data processing, typically based on either the batch [5, 47, 11] or stream [28, 9, 23, 2] processing models. These systems allow users to express a dataflow of logical operators (Figure 1), while the system automatically handles data distribution, task scheduling, and fault tolerance. Data processing in ML pipelines often consists of pure map transforms and thus can easily be expressed with this API [30]. However, these systems were designed for homogeneous CPU-only clusters. Heterogeneous ML pipelines have two key properties that limit performance and scalability when executed on CPU-centric data processing frameworks.
(a) Batch inference.
(b) Training Stable Diffusion.
Figure 1: Heterogeneous applications in batch inference and training, represented as logical dataflow graphs (nodes are operators). (a) represents a typical pipeline for video or image generation. (b) represents a pipeline for training stable diffusion model. In (b), UNet model is replicated for data-parallel training.
First, different operators require different degrees of physical parallelism, i.e. the number of concurrently executing instances. Thus, a key system requirement is to decouple the physical parallelism of each operator. For example, one may use many CPU threads to download an input dataset from cloud storage to match the throughput of a single GPU. This requires buffering intermediate data in memory until the downstream operator is ready to consume them. As data modalities such as text, images, and video proliferate [32, 40, 16, 38], the intermediate data size can be significant and unpredictable. Meanwhile, memory is more challenging to multiplex, as oversubscription is costly. Maintaining throughput while staying under memory limit is critical.
Second, some resources are more valuable than others. Thus, it is critical to isolate failure domains to the failed resource. For example, GPUs are more expensive than CPUs, but CPU failures are more likely because they are deployed in higher numbers and more often on spot instances [46]. Ideally, a CPU failure should have little impact on GPU execution. This is challenging to achieve simultaneously with exactly-once record processing and minimal run-time overheads. Even with stateless transforms, the system must track which records have been processed for each operator to avoid duplicating or dropping records during failover.
Thus, a distributed data processing system for ML pipelines must: (1) dynamically adjust each operator’s parallelism according to actual compute and memory usage, while (2) minimizing run-time and recovery overheads from failures and dynamic re-scaling. Current batch and stream processing systems can achieve one but not both. Batch processing systems execute operators in synchronous stages of stateless and immutable tasks, replaying them upon failure. This method, known as lineage reconstruction, offers high fault tolerance and elasticity but limited support for dynamic parallelism [11, 47]. Stream processing systems execute operators with asynchronous stateful executors that can decide locally when to materialize records. However, this decentralized approach requires a recovery method that imposes either high run-time overhead from logging [2, 28] or high recovery overhead from global checkpointing and rollback [9, 28].
To address these challenges, we present the streaming batch model, a hybrid model for efficient and fault-tolerant heterogeneous execution. Similar to the batch processing model, a centralized scheduler partitions data across a stage of stateless tasks. Tasks can run on any executor and are recovered via lineage reconstruction, enabling elasticity and fault tolerance while avoiding unnecessarily expensive logging. However, similar to the stream processing model, stages are executed asynchronously, allowing data to be streamed to the next operator and dynamically repartitioned.
(a) Batch processing model.
(b) Stream processing model.
(c) Streaming batch model (Ray Data).
[…]
We present Ray Data, a distributed streaming batch system for heterogeneous workloads such as batch inference and ML training. Ray Data implements a logical dataflow API with an ahead-of-time execution planner and a run-time scheduler. The execution planner decides the initial number of partitions per operator. During execution, operators decide when to materialize a task’s intermediate partitions, allowing dynamic repartitioning based on local memory usage. Meanwhile, the centralized scheduler maintains a global view of running tasks and materialized partitions and can accordingly adjust the physical parallelism of each operator, i.e. its memory and compute allocation, while enforcing overall memory limits.
Ray Data uses Ray [27] as a distributed task backend. We extend Ray’s lineage-based recovery with a key feature needed for streaming batch execution: dynamic repartitioning.
We evaluate Ray Data on batch inference and ML training workloads that span diverse resource requirements (CPUs vs. GPUs), storage (local disk vs. cloud), and modalities (image vs. video). Ray Data outperforms batch and stream processing systems such as Spark and Flink with 3–8$\times$ better throughput. Ray Data is also able to match the throughput of single-node ML-specific data loaders such as tf.data and PyTorch DataLoader while additionally leveraging distributed and heterogeneous clusters. On a Stable Diffusion training benchmark, Ray Data can improve training time by 31% by leveraging a pool of 72 heterogeneous GPUs. In summary, we contribute:
  • The streaming batch model, an efficient and fault-tolerant execution model for distributed heterogeneous processing.
  • An online and heterogeneity-aware scheduling policy for streaming batch systems that can enforce total memory limits and maximize total compute utilization.
  • Ray Data, a fault-tolerant, memory-efficient, and autotuning implementation of the streaming batch model.
2 Background
We overview key aspects and limitations of the batch and stream processing models, including ML data loaders, by analyzing how effectively each system can:
  • Manage memory for intermediate data between operators.
  • Maximize utilization of heterogeneous compute resources.
  • Minimize overheads for cluster failures and re-scaling.
2.1 Applications
We target ML training and inference pipelines that require data pre- and post-processing using CPUs, GPUs, or both. Our goal is maximizing throughput, while providing sub-second latency when used as data loaders for ML training [30]. Similar to current ML dataloaders [30, 35], we primarily target map-style per-row transforms. Operations that require all-to-all shuffle exchanges, such as sort and group-by, are also supported in the system but are already discussed in [25].
Figure 1(a) shows a typical batch inference pipeline, which uses CPUs to load and filter a dataset from cloud storage (e.g., HDFS or S3), GPUs to produce predictions, then CPUs to upload the results. A key characteristic is that the different operators may require different degrees of parallelism. For example, load may require many CPU threads to download an input dataset from cloud storage, while filter may only require one per core. Also, CPUs typically vastly outnumber GPUs in a cluster, so predict runs at much lower parallelism.
Figure 1(b) shows a typical distributed training pipeline for the Stable Diffusion model [38]. The pipeline uses CPUs to load and preprocess image-text pairs, GPUs to produce encodings with a pre-trained Encoder model, then GPUs to train a UNet model. Even GPU operators can leverage heterogeneity: colocating Encoder and UNet on the same GPUs can lead to lower training throughput, as it takes valuable resources from the UNet. If the operators’ physical resources are decoupled, then Encoder can be run on lower-end GPUs. This reduces overall cost by increasing utilization on the UNet GPUs.
2.2 Batch Processing Model
Batch processing systems are designed to allow a task to run and therefore recover on any executor. The system transforms the user-provided logical DAG (Figure 1(a)) into a physical DAG of stages of data-parallel tasks (Figure 2(a)). Tasks are stateless and materialize their input and output partition(s). This is key to lineage-based recovery, which logs only the logical DAG and re-executes tasks to recover lost partitions. Elastic scaling is supported by simply adding or removing executors. Examples include MapReduce [11], Apache Hadoop [5], Apache Spark [47], Spark Streaming [48], and Apache Flink in BATCH execution mode [9].
However, to make this recovery method practical, the system imposes two significant restrictions on execution. First, each stage must fully execute and materialize its outputs before executing the next stage. This simplifies scheduling and recovery, as a (re)scheduled task never idles waiting for its inputs. To reduce overheads from materialization, consecutive map operators are often fused into one stage [47]. This is effective when operators require the same resources. For example, suppose the load and filter operators in Figure 1(a) are fused. If each uses 1 CPU, then this has no impact on compute utilization, but greatly reduces the memory footprint, as data can be loaded and filtered one batch at a time.
Unfortunately, when different stages require different resources, operator fusion can cause significant underutilization. For example, with 4 CPUs and 2 GPUs, if Figure 1(a) fused all operators, then we could execute at most two tasks at a time. This would underutilize the CPUs while leaving the GPUs idle during CPU execution. Disabling fusion of heterogeneous operators, as in Figure 2(a), is also imperfect, as it requires materializing each stage’s outputs. This prevents pipelining between CPUs and GPUs, while producing high memory pressure and likely disk spilling.
Second, the data partitioning must be determined before execution so that it can be recorded in the lineage. This prevents the system from using run-time information such as the in-memory size of intermediate data rows when deciding the partitioning strategy. Thus, even if pipelined stage execution were supported, there can still be high memory pressure from individual intermediate partitions that are too large. For example, suppose that Figure 1(a) were run on a dataset of 100 videos, each 10MB on disk but 20GB decoded [7]. The system may choose 100 partitions to ensure good load-balancing for load, but each task would require 20GB memory! While some control is exposed to the user, typically one can only specify a target number [47], again before execution. Thus, it is left to the user to predict each operator’s memory usage, then manually configure the number of partitions.
2.3 Stream Processing Model
Stream processing architectures optimize for online scenarios and asynchronous execution. Examples include Naiad [28], Apache Flink [9], Spark Continuous Processing [42], MillWheel [2], and Apache Kafka [23]. Typically, each logical operator is assigned a parallelism, by the user or the system, which determines how many physical operator instances to create. For example, the parallelism of A-B-C in Figure 2(b) is 4-2-4. Consecutive logical operators that have the same resource requirements and parallelism are fused to avoid materialization. Physical operators execute asynchronously on their input stream(s) and exchange record batches directly, without involving a centralized scheduler.
Each physical operator processes a pre-determined stream partition. For example, in Figure 2(b), A8 is pre-assigned to CPU3. In contrast, Figure 2(a) executes A8 earlier, on any free CPU. However, unlike batch processing systems, physical operators are stateful and can choose at run time how many records to process at a time. For example, in Figure 2(b), the system assigns CPU3 1/4 of the total key range for operators A and C, but the executor chooses when to materialize A4 vs. A8 and C4 vs. C8. Typically, the executor will accumulate records up to a time or memory limit, then materialize and send the batch to a downstream executor. If the downstream executor is overloaded, backpressure is applied to limit memory.
Physical operators are tied to an executor and assigned one global partition, so reconfiguration is important for load-balancing. Reconfiguration efficiency is a common challenge [43], and especially so for heterogeneous pipelines. First, for heterogeneous operators, fusion is often impractical. For example, Figure 2(b) does not fuse A and C because B requires a GPU. Determining the optimal parallelism for A and C can be challenging without run-time information. Second, heterogeneous clusters can greatly reduce cost, but efficient elastic scaling and failure handling are critical.
There is a common tradeoff among stream processing systems between run-time overheads vs. reconfigurability and recovery overheads. Many systems use asynchronous global checkpointing [28, 9]. This typically imposes low execution overheads, but any failure causes a global rollback to the last checkpoint. Also, global checkpoints must coordinate each process’s local checkpoints, so static membership is often assumed for simplicity [13]. Thus, reconfiguration typically introduces long pauses, as it requires taking then restarting from a global checkpoint with the new configuration [9].
Other systems use logging [2, 23]. This allows for fast recovery via log replay and fast repartitioning of physical operators. However, intermediate records must be durably logged before releasing to the downstream operator, adding higher run-time overheads. This is acceptable for online systems that interact frequently with the external world, but ML pipelines are typically offline systems where most intermediate values are safe to rollback and thus do not need durability.
ML data loaders.
ML-specific data loaders are single-node systems that maximize I/O and CPU bandwidth to improve local GPU utilization. Examples include tf.data [30] and PyTorch DataLoader [35]. These systems can be viewed as a special case of stream processing: they launch a fixed pool of worker threads or processes at run time, which continuously load, preprocess, and feed data to an end GPU consumer. The pool must be tuned to match the GPU’s throughput without running out of memory. tf.data automatically adjusts operator parallelism at run time to maximize GPU utilization [30].
However, in general these data loaders are narrow in scope: (1) they do not support distributed execution, and (2) GPUs are always assumed to be sinks. (1) prevents load-balancing and heterogeneous clusters. Both (1) and (2) make it impossible to leverage heterogeneous GPUs in the Stable Diffusion example in Figure 1(b), as well as batch inference pipelines that alternate CPU and GPU operators. Adapting data loaders to support these applications would require effort equivalent to re-building a distributed data processing framework.
Batch [11, 5, 47] Stream [28, 9, 2, 23] PyTorch DL [35] tf.data [30] Streaming batch (Ray Data)
Automatic partitioning $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Dynamic repartitioning $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$: streaming repartition
Min. pipeline granularity Stage Partition Partition Partition Partition
Dynamic parallelism $\checkmark$ $\checkmark$/$\times$ $\times$ $\checkmark$ $\checkmark$
Distributed execution $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
Fault tolerance method Lineage Logging/Checkpointing None Checkpointing Lineage
Min. rollback granularity Partition Record/Epoch Epoch Epoch Partition
Table 1: Features for heterogeneous and distributed processing. Dynamic repartitioning is important for memory efficiency. Finer pipeline granularity improves utilization for heterogeneous resources. Dynamic parallelism is important for adaptivity. Stream processing systems use logging for dynamic parallelism with zero downtime and record-level rollback, or checkpointing, which requires a checkpoint to reconfigure and global rollback on failure (§ 2.3).
3 Overview: The Streaming Batch Model
The streaming batch model (Figure 2(c)) uses task-based execution with a centralized scheduler. Tasks can run on any executor, similar to batch processing (Figure 2(a)). However, tasks across different stages are pipelined and dynamically repartitioned (Figure 3(b)), similar to stream processing systems (Figure 2(b)). Table 1 shows a full feature comparison.
There are two primary challenges: (1) extending lineage reconstruction to support streaming batch execution, and (2) building a centralized scheduler that adaptively manages all running tasks and intermediate partitions.
Challenge 1: Fault-tolerant, memory-efficient partitioning.
Batch processing tasks are often memory-inefficient when operators require different parallelisms. For example, Figure 3(a) shows a two-stage pipeline where parallelism=1 and 3 for stages A and B, respectively. Repartitioning A1 requires a stage barrier, causing high memory usage.
Our goal is to achieve the memory efficiency of stream processing but maintain the low run-time and recovery overheads of lineage-based recovery. In particular, we want to allow the user to set a single target partition size, instead of requiring the user to estimate the number of partitions per operator. This is challenging because then the number of partitions is unknown until run time. Meanwhile, lineage-based systems typically require logging the operations before execution.
To address this, we propose a streaming repartition (§ 4.2.1): a task may dynamically generate multiple output partitions, based on its real-time memory consumption. For example, in Figure 3(b), if A1 experiences memory pressure, it can output a partition early. As we also support pipelined stage execution, this reduces peak memory usage: B tasks can begin and release their input partitions while A1 is executing.
For recovery, we extend Ray’s lineage-based recovery [44] (§ 4.2.2). Ray includes a distributed object store, which we use for intermediate partitions, and task-parallel execution. Like other lineage-based systems, Ray requires immutable task definitions and does not allow task outputs to be read before task completion. To support streaming repartition, we add support in Ray for tasks with dynamic and streaming outputs.
(a) Repartitioning in batch processing.
(b) Streaming repartition.
Figure 3: Streaming repartition (b) allows executors to: (1) locally decide when to output partitions, and (2) pipeline execution with the next operator. Overall, this reduces peak memory usage compared to repartitioning in batch processing systems (a).
(a) Pessimistic scheduling.
(b) Optimistic scheduling.
Figure 4: Scheduling under memory pressure. Green represents CPU executors’ local memory capacity (1 partition per CPU, 2 total). Pink represents the system’s shared memory capacity for intermediate data (1 partition total). (1): If shared memory is full, executors must stall and buffer outputs locally until space is made, either via spilling to disk or executing downstream tasks. (b) is faster because it schedules A3 as soon as possible but doesn’t stall CPU0.
Challenge 2: Dynamic reconfiguration.
One advantage of the streaming batch model is that it uses a centralized scheduler to schedule all tasks, giving the scheduler a global view of all resources. This enables enforcing a hard limit on memory used by intermediate data partitions. However, this must be done without unnecessarily stalling compute tasks. Memory is shared among all operators, and oversubscribing memory can cause extreme slowdowns from stalls or spilling to disk. When memory is limited, the choice of which task to execute and when is not obvious.
For example, consider Figure 4(a). Each CPU executor has local memory capacity (green) for one partition, and the system has shared memory capacity (pink) for one partition. In phase (1), A2 must stall and buffer its outputs locally until B1 completes. A conservative scheduler additionally waits for phase (2) before scheduling A3, to avoid stalling CPU0. Scheduling A3 optimistically so that it finishes simultaneously with B2, reduces overall run time (Figure 4(b)). Applying such optimizations requires dynamic profiling and reconfiguration.
To address this challenge, we introduce an adaptive scheduler in Section 4.3. The key idea is to fairly allocate shared compute resources between operators, while estimating future memory availability from run-time information to enable optimistic scheduling.
Method Description
read Read items from files.
map Transform each item.
map_batches Transform a batch of items. Useful for controlling GPU batch size.
flat_map Transform each item and flatten the results.
filter Return items that match a predicate.
limit Truncate to the first N items.
write Write items to files.
iter Return an iterator of items.
iter_split Split into N iterators.
materialize Materialize all items.
Table 2: A subset of the Ray Data Dataset API. The bottom four are consumption APIs that trigger execution, while the others are lazy.
3.1 The Dataset API
A Dataset represents an application pipeline. Datasets are lazily created, by reading files or applying transforms to an existing Dataset (Table 2). A Dataset is materialized through a write operation, e.g., to cloud storage, or by iterating over the items in memory.
A key part of the API is the ability to express resource requirements. Resource requirements are a map from resource name to float value and may be passed as an option to the transforms. By default, each transform requires 1 CPU. Resource names can be CPU, GPU, or a custom resource label.
Most of the map-style transforms take a stateless and pure user-defined function (UDF) as an argument. For operations that require significant initialization time, such as a model loaded into GPU memory, we also support stateful UDFs that can be instantiated once and called multiple times on different items. We assume that all UDFs are pure, to enable lineage-based recovery.
Figure 5: Ray Data architecture overview. Ray Data executes as a Ray library. The Ray Data scheduler executes as a Ray driver, dispatching tasks to Ray workers (dashed arrows).
3.2 Executing a Ray Data Program
After creating a Dataset, the user triggers execution by calling one of the consumption APIs, as seen in (1) in Figure 5. The Dataset is represented as a DAG of logical operators, as in Figure 1. The system’s query planner then compiles this logical DAG into a DAG of physical operators ((2) in Figure 5). The query planner applies operator fusion and decides the number of partitions to use for the first operator (§ 4.1). Each physical operator defines a transform to apply to each row and metadata such as the resources required. Physical operators are decoupled from executors, and their parallelism may not be determined until run time.
Ray Data uses Ray as a task backend, storing intermediate data partitions in Ray’s distributed object store. During execution, the Ray Data scheduler dispatches each physical operator as one or more Ray tasks ((3) in Figure 5). Stateless UDFs are executed as Ray tasks, which are Python processes that can run anywhere in the cluster. Stateful UDFs are executed as tasks running on Ray actors, which are long-running Python processes that preserve state. Actors acquire resources for their lifetime. Thus, if multiple stateful UDFs require the same resource, multiplexing is possible by sharing an actor pool between UDFs.
For each physical operator, the Ray Data scheduler keeps a queue of ready partitions, either a Dataset.read input, e.g., a batch of input filenames, or an intermediate partition. We use Ray as a decentralized dataplane, so that the Ray Data scheduler only needs to keep references to partitions, not the physical data [44]. In a loop, the scheduler scans the current resource availability, chooses a ready partition, then passes it by reference to a new Ray task along with a description of the physical operator to execute. We describe the policy in Section 4.3.
We build upon and extend Ray’s fault tolerance. Ray provides lineage-based recovery and stores objects in separate processes from the workers, so that individual executor failures do not impact materialized partitions [44]. If a materialized partition is lost due to node failure, Ray automatically recreates it by resubmitting the tasks in its lineage. In Section 4.2.1, we describe the extensions we made to Ray’s recovery method to support streaming repartition.
4 System Design
4.1 Query Planning
The query planner takes as input a logical DAG and parameters such as the target maximum partition size. It generates a physical DAG, then applies a series of optimizations on the physical DAG, for example, adjacent operators that require the same resources are fused into one physical operator.
During execution, tasks may dynamically repartition the data. The initial number of partitions to use for a read operator is decided by the query planner. It needs to be sufficiently large to utilize all available execution slots (usually CPUs), but not so large that each partition is tiny, which can increase system overheads. It is also upper-bounded by the number of input files. By default, we aim to produce partitions that are 1—128 MB in size. We compute the initial number of partitions based on the following heuristics: the number of initial execution slots, the estimated output size of the read operator, and the user-requested value, if any.
Some Dataset consumption APIs (Table 2) induce transformations on the physical DAG. The Dataset.write call is appended to the DAG as a map that writes items to external storage. The iter API returns a stream of output records. Under the hood, this is implemented by fetching and buffering output partitions. The iter_split call shards the outputs into N streams, each of which can be passed to a different process. This is useful for cases such as distributed data-parallel training where the dataset is sharded among N trainers. To implement iter_split, the query planner launches a Ray actor before execution to coordinate the dynamic partition assignment. The stream readers fetch and buffer output partitions from this coordinator actor. Partitions are passed by reference to avoid coordinator bottleneck.
4.2 Query Execution
During execution, the Ray Data centralized scheduler has a global view of the executing tasks and the available resources. The scheduler repeatedly executes the following loop:
  • Wait for an executing task to materialize an output partition. Tasks may produce multiple output partitions; if this was the last, then mark the task’s resources as free.
  • While there are free resources and ready partitions, launch new tasks using the policy described in Section 4.3. Mark the task’s required resources as used.
The scheduler passes in the task description: (1) a closure of the physical operator to execute, (2) references to the task’s input partition(s), and (3) the target output partition size. We leverage Ray to ensure that the task’s input partition(s) are made local to its executor.
4.2.1 Streaming repartition
Although the query planner makes an initial estimate of the number of output partitions to use, this value may not be optimal. When a physical operator’s outputs are much larger or smaller than its inputs, the initial partitioning will produce too-large or too-small inputs. Too-large partitions can cause out-of-memory failures from buffering many records at the executor. On the other hand, too-small partitions are also inefficient; while the Ray Data scheduler only stores references to partitions, it still must manage some metadata per partition and carry out RPCs to schedule corresponding task(s).
Ray Data introduces a technique called streaming repartition to handle such cases (Figure 3(b)). To support this, we extend Ray with remote generator tasks, which enable Ray tasks to produce a dynamic number of outputs, and to pipeline execution with the task’s caller (the Ray Data scheduler). Whenever the task produces a new output, it notifies the Ray Data scheduler via RPC. Upon receipt, the scheduler can immediately launch a downstream task. Meanwhile, the upstream task can continue producing its next output partition.
Ray Data tasks take a target partition size from the scheduler and determine locally how to partition their outputs. When executing a task, the worker accumulates processed rows to a local output buffer. Once the output buffer exceeds the maximum target partition size (128 MB by default), it uses the yield keyword in Python to serialize and materialize the partition in Ray’s object store. If a logical operator produces much less data than it consumes, a task may produce a too-small partition. To handle this, we simply coalesce partitions by having the Ray Data scheduler pass multiple partitions from different upstream tasks to a single downstream task.
4.2.2 Failure recovery
Ray provides automatic recovery for objects (intermediate partitions) as long as (1) the driver is alive, (2) the tasks that created them are deterministic and side effect-free [44], and (3) the task arguments and outputs are immutable. For generator tasks, (3) is no longer true, because we do not know at submission time how many outputs the task will produce. However, we note that if (2) is true, then streaming repartition can be made deterministic. In particular, given a target partition size, a pure transform, we ensure that a Ray Data task will produce the same stream of output partitions if executed on the same input partition(s).
To support failure recovery for generator tasks, we modify Ray’s recovery subsystem to handle tasks with an unknown number of outputs. Generator tasks are initially launched with an unknown number of outputs. On the first successful execution, the task’s caller records the number of outputs that the task produces. If any of the task’s outputs are lost, we recover by re-executing the entire task. If the task produces a different number of outputs, we throw an error.
Similar to other batch and stream processing systems [9, 34, 47, 30], if the centralized scheduler dies, Ray garbage-collects the job and it must be re-executed from the beginning. In the future, this case can be optimized through known techniques for asynchronous global checkpointing [2, 9].
4.3 Adaptive Scheduler
The goal of the scheduler is to minimize the job completion time while keeping the total memory usage of intermediate data below the system limit. To achieve this, the scheduler uses a principle of equalizing the processing rates of each operator (in bytes per second). Intuitively, if processing rates are not equal, slower operators will accumulate pending inputs, eventually exhausting the memory buffer. Each operator’s processing rate can be controlled by deciding how many tasks to run in parallel for that operator. The processing rates are estimated online using run-time statistics, including operator task durations and average task input–output data size ratios, because these properties are difficult to predict ahead of time, and could vary depending on the actual data being processed.
4.3.1 Algorithm
The input to the scheduler is as follows:
  • The physical DAG of operators (§ 4.1). Each operator can process data items in parallel tasks and is annotated with its resource requirements, e.g., {GPU:1}.
  • Resource limits: the total number of CPU, GPU, or custom resource slots, and $totalMemoryCapacity$ of the system.
Algorithm 1 Adaptive Scheduler
1:Initialize $budget\leftarrow totalMemoryCapacity$
2:while not all operators are done do
3:     Update resource utilization and run-time estimates
4:     Update budget $\triangleright$ Algorithm 2
5:     if $budget\geq\mathrm{outputPartitionSize}(source)$ then
6:         Launch task of $source$
7:         $budget\leftarrow budget-\mathrm{outputPartitionSize}(source)$
8:     end if
9:     $Q\leftarrow\emptyset$triangleright Set of qualified operators
10:     for each operator $op$ in DAG do
11:         if hasInputData($op$) and
12:     hasAvailableResources($op$) and
13:     hasOutputBufferSpace($op$then
14:              $Q\leftarrow Q\cup\{op\}$
15:         end if
16:     end for
17:     if $Q\neq\emptyset$ then
18:         $selected\leftarrow\operatorname*{arg\,min}_{op\in Q}\text{bufferedOutputsSize}(op)$
19:         Launch task of $selected$
20:     end if
21:end while
Algorithm 1 describes the scheduling loop in step (3) of Figure 5. From all qualifying operators, it picks the one with the least amount of data accumulated in its output buffer. The intuition is that this operator would be producing data at a slower rate than it is consumed, and thus needs more parallelism. This policy works well for equalizing the operator processing rates when there is enough memory to store the intermediate data while all executors are utilized.
However, when memory is constrained, the policy (without lines 4–8) will result in resource under-utilization, as seen in Figure 4(a). The ideal solution is to keep the pipeline full and start source tasks, i.e. tasks for the source operator, as early as possible, as in Figure 4(b). To achieve this, lines 4–8 add a higher-priority optimistic policy for scheduling source tasks, described further in the next section.
4.3.2 Input Rate Control
The input rate, i.e. the rate at which source tasks are scheduled, must approximate the pipeline’s overall throughput: If the source tasks are launched too slowly, the downstream operators will have no input to process, wasting compute resources. Conversely, if the source tasks are launched too aggressively, then these tasks may starve downstream operators and eventually cause slowdowns from back-pressuring of the source operator and/or spilling to disk.
We use a dynamic memory budget algorithm to regulate the rate at which source tasks [^1] For DAGs with multiple sources, the launch rate for each operator should be proportional to their data output size. are launched. Intuitively, the budget is an optimistic estimate of the memory available for new data partitions to enter the system. When a source task is launched, we deduct its estimated output size from the budget. At every second, the budget is updated using Algorithm 2, which estimates the rate at which data leaves the pipeline.
Algorithm 2 Memory Budget Update (runs every second)
1:$P\leftarrow 0$triangleright Total processing time per partition
2:$\alpha_{0}\leftarrow 1$triangleright $\alpha_{i}:=$ Input:Output size ratio for $op_{i}$
3:for $i\leftarrow 1$ to $numOps$ do
4:     $E_{i}\leftarrow\text{availableExecutionSlots}(op_{i})$
5:     $T_{i}\leftarrow\text{estimatedTaskDuration}(op_{i})$
6:     if $op_{i}$ is not $source$ then
7:         $I_{i}\leftarrow\text{estimatedInputSize}(op_{i})$
8:         $O_{i}\leftarrow\text{estimatedOutputSize}(op_{i})$
9:         $\alpha_{i}\leftarrow\alpha_{i-1}\cdot O_{i}/I_{i}$
10:     end if
11:     $P_{i}\leftarrow(T_{i}/E_{i})\cdot\alpha_{i-1}$
12:     $P\leftarrow P+P_{i}$
13:end for
14:$budget\leftarrow budget+\mathrm{outputPartitionSize}(source)/P$
We will walk through the algorithm using the following example: load (CPU) $\to$ transform (CPU) $\to$ inference (GPU) pipeline, running on a cluster with 8 CPUs and 4 GPUs.
  • Consider the first non-source operator: transform. Assume that the number of available execution slots to run the task is $E_{1}=6$ (out of 8 CPU slots). Assume the average task duration is $T_{1}=12$ seconds. Then the processing time of this stage is $P_{1}=T_{1}/E_{1}\cdot\alpha_{0}=12/6\times 1=2$, where $\alpha_{0}$, the output multiplier, is initialized to 1. In other words, transform takes 2s to process a source partition on average.
  • Assume the transform’s average output is double the size of its input, i.e. $\alpha_{1}=2$. Now consider inference. Assume the number of available execution slots is $E_{2}=4$ (GPU), and the average task duration is $T_{2}=2$ seconds. Then $P_{2}=T_{2}/E_{2}\cdot\alpha_{1}=2/4\times 2=1$, i.e. the inference operator takes 1 second per source partition.
  • Adding them up, $P=2+1=3$ seconds per source partition. In other words, every ~3s, the budget will be replenished to allow for one more source task to run.
If the run-time estimates of each operator’s processing rates are perfectly accurate, i.e. if there is no variance in the processing rates, it can be shown that the schedule produced is optimal. However, when there is variance in processing times or output sizes, the budget algorithm could overestimate the overall processing rate. Nevertheless, the algorithm is stable because it creates a negative feedback loop. If it overestimates the pipeline processing rate, more source tasks might launch and temporarily cause backpressure, or objects in the buffer to spill to disk. However, since these tasks still occupy execution slots, they will reduce the parallelism of downstream operators and lower the replenishment rate of the budget, which in turn limits the source task launch rate. This will in turn release more resources for downstream operators to run, consuming the buffered intermediate data, and bringing the pipeline throughput back to equilibrium.
Ray Data also provides a conservative scheduling policy in which a task is launched only when its output space in the memory can be guaranteed, similar to Figure 4(a). This policy enforces a hard memory limit and never spills, albeit at the risk of under-utilizing of executor slots when memory is limited.
4.4 Implementation
Current batch and stream processing systems could in principle be modified to use the streaming batch model, but would require fundamental changes to their execution models. For example, Spark would need to support pipelined stage execution, while Flink would need to support zero-downtime reconfiguration. We choose to implement Ray Data on top of Ray because Ray exposes a lower-level execution model based on dynamic task execution. Ray provides powerful features such as automatic data movement, lineage-based recovery [44], and automatic disk spilling [25]. This makes it convenient to build centralized schedulers like Ray Data while leveraging Ray as a decentralized dataplane. However, Ray also treats the task logic, inputs, and outputs as black boxes. Thus, it is difficult to directly extend Ray with data processing-specific features such as dataset partitioning, streaming repartition, and pipeline-aware task scheduling. Instead, we implement Ray Data as a Ray library, which allows us to build such features with minimal changes to the Ray core. Ray Data is written in ~90k Python LoC, including ~45k LoC for the query planner, scheduler and executor logic.
5 Evaluation
We evaluate a range of heterogeneous workloads with a focus on multimodal batch inference and training. We use benchmarks taken from the MLPerf suite [26, 36], and supplement with additional image-to-image, video-to-video, and video classification benchmarks. We aim to answer:
  • § 5.1: How does the streaming batch model compare to traditional batch or stream processing models when running heterogeneous ML workloads, in terms of throughput and adaptivity vs. fault tolerance?
  • § 5.2: How does distributed and heterogeneous execution improve on throughput per dollar compared to ML-specific single-node data loaders?
  • § 5.3: How well do all systems adapt to memory pressure in heterogeneous settings, and what are the system overheads?
We compare the following systems:
  • Batch processing (Figure 2(a)): Apache Spark 3.5.1.
  • Stream processing (Figure 2(b)): Apache Flink 1.19.0.
  • Single-node stream processing: tf.data [30]. A data loader for ML training. tf.data uses multithreading and can reconfigure the number of threads per operator.
  • Single-node stream processing: PyTorch DataLoader (PyTorch DL) [35]. Similar to above, but uses multiprocessing and requires manual fusing of all operators.
  • Streaming batch: Ray Data (implemented over Ray 2.40.0). We also modify Ray Data to emulate batch processing (Ray Data-staged) and stream processing (Ray Data-static), for apples-to-apples comparison of the execution models in § 2. Ray Data-staged materializes each stage before starting the next, while Ray Data-static sets a static parallelism per operator and disables the adaptive scheduler described in § 4.3.
For training workloads, we compare only against tf.data and PyTorch DataLoader because these systems were custom-built for data preprocessing for training. They do not support distributed execution and target CPU-only preprocessing for a co-located GPU trainer, making it difficult to run batch inference, which alternates between CPU and GPU stages. Thus, we use Spark and Flink as comparisons for batch inference.
(a) Image generation.
(b) Video generation.
(c) Fault tolerance.
(d) Scalability.
Figure 6: (a) Image generation throughput comparison. Ray Data-dynamic (blue) can match the performance of hand-tuned Ray Data-static (red) with static parallelism, and outperform other systems. Ray Data-staged execute synchronously; no results are available until the last stage begins at ~330s. (b) Video generation throughput comparison. Ray Data-dynamic (blue) achieves 28% better throughput compared to Ray Data-static (red) with static parallelism. (c) Fault tolerance comparison. Ray Data can handle isolated executor failures (blue) with negligible impact on throughput, and node failure (red) without restarting the job. The emulated checkpoint-and-restore (pink) leads to job downtime and longer completion time due to recomputation. (d) Scalability analysis. Ray Data is able to achieve strong scaling with respect to the number of nodes.
5.1 Inference: Comparison to distributed batch and stream processing
5.1.1 Image-to-Image Generation
Image-to-image generation uses a generative model to produce a new image from a source image and prompt. Section 5.1.1 consists of the steps: (1) read_images: download the InstructPix2Pix [8] image dataset from S3, (2) decode: decode JPEG images to tensors, (3) preprocess: normalize tensors, (4) Img2ImgModel: use Stable Diffusion [38] to generate a new image based on the input image and prompt, and (5) encode_and_upload: Encode to JPEG and upload to S3.
{listing}
[ht]
ray.data.read_images(INPUT_PATH).map(decode).map(preprocess)
.map_batches(Img2ImgModel, batch_size=B, num_gpus=1)
.map_batches(encode_and_upload, batch_size=B)
Image-to-image generation expressed in Ray Data API.
Figure 6(a) shows throughput over time of the different execution models on 1 g5.2xlarge VM (8 vCPU, 1 A10G GPU). The *-fused baselines fuse all operators. This avoids materialization but limits overall parallelism to the scarcest resource, in this case 1 GPU. The resulting throughput is thus the lowest, confirming that operator fusion is undesirable on heterogeneous resources. Spark-staged and Ray Data-staged disable fusion of CPU and GPU operators to decouple their parallelisms. Stages execute synchronously, so no results are available until the last stage begins at ~330s. The overall throughput is also significantly lower because all intermediate stage results need to be materialized and in this case spilled to disk.
For Flink, we also disable fusion, by setting parallelism to 8 and 1 for CPU and GPU operators, respectively. Flink uses multithreading to share each CPU among its multiple operators (read+decode+preprocess vs. encode_and_upload). Execution is pipelined across physical operators, and only a fraction of intermediate records are materialized at once, so Flink can produce results almost immediately. However, there is significant overhead due to serialization of image data between the Python UDF and the Java-based Flink. Thus, Flink’s throughput is 70% lower than Ray Data’s.
Because of Flink’s serialization overheads, we use Ray Data-static to emulate a stream processing baseline. We manually determine the best static parallelism per operator before execution. This achieves the best throughput, at 4 images/s. Ray Data-dynamic is the default system, with the adaptive scheduler (§ 4.3). This automatically converges to the same throughput as Ray Data-static.
Takeaways:
(1) Stream and streaming batch models outperform batch models for heterogeneous pipelines, due to asynchronous stages and lower peak memory footprint, (2) Ray Data’s adaptive scheduler matches the best stream processing baseline (Ray Data-static), with no configuration needed.
Figure 7: Training ResNet-50.
Resources Images/s Run time (hours) Total cost
PyTorch DL (stream) 4×4$\times$4 × p4de.24xlarge 2,811 111.3 \$18,192
Ray Data-staged (batch) 4×4$\times$4 × p4de.24xlarge 0, then 4,068 90.3 (-19%) \$14,753 (-19%)
Ray Data (streaming batch) 4×4$\times$4 × p4de.24xlarge 4,075 76.8 (-31%) \$16,275 (-11%)
40×40$\times$40 × g5.2xlarge
Figure 8: Run time and cost for one epoch of Stable Diffusion pre-training.
5.1.2 Video-to-Video Generation
Video-to-video generation uses a generative model for use cases including super-resolution, interpolation, or de-blurring [10, 20, 45]. Compared to § 5.1.1, video generation adds significant memory pressure: each video decodes to many frames. Also, high workload variability is common due to varying video lengths, resolutions, and encodings. Thus, we use this workload to evaluate the system’s ability to adapt compute and memory allocations based on real-time usage.
We run a 3-stage pipeline on a g5.2xlarge VM: (1) download+decode (CPU): Download the YouTube-8M dataset [1] from S3. Video resolution ranges from 320p to 720p, with lengths 2–5 minutes, encoded using H.264. Each video is decoded into multiple 128-frame sequences, the maximum batch size allowed by GPU memory. Earlier videos are low-resolution. (2) generate (GPU): Run the RealBasicVSR [10] super-resolution model to produce HD frames. (3) encode+upload (CPU): Encode the HD frames using H.264, then upload the videos to S3.
Figure 6(b) compares the throughput over time of Ray Data-static and Ray Data-dynamic, the two best systems in § 5.1.1. Ray Data-static emulates stream processing and allocates the pre- and post-processing stages of 4 CPUs each. Throughput is initially high but later drops because this initial configuration is not optimal for the later high-resolution videos, which are more memory-intensive and take longer to download and decode. Meanwhile, with Ray Data-dynamic, the Ray Data scheduler dynamically re-balances tasks according to task duration and memory usage, achieving 28% better throughput.
Takeaway:
Ray Data adapts quickly to changing workloads because tasks can run on any executor and the adaptive scheduler can dynamically reallocate resources.
5.1.3 Fault tolerance in heterogeneous clusters
We first demonstrate Ray Data’s ability to scale with heterogeneous clusters. We run the VideoMAE [41] for video classification, on the Kinetics-700-2020 dataset with 635,000 videos and 871 GB in size. This workload is similar to that in § 5.1.2 except with less postprocessing. Data is stored on Amazon S3. Figure 6(c) shows the throughput over time when processing 10% of the dataset on 1 g5.xlarge node (4 vCPU, 1 GPU). The single-node configuration produces 6.2 videos/s, and is bottlenecked by CPU preprocessing. Ray Data can scale data preprocessing independent of GPU inference. We launch a m7i.2xlarge node (8 vCPU) to create a heterogeneous Ray Data cluster with 12 vCPUs and 1 GPU. Overall throughput then reaches 21.7 videos/s, or 94% of the maximum GPU throughput. The overall job cost is also 60% less, as CPU-only nodes are more cost-efficient. We further test scaling the cluster up to 16 nodes, or 8 GPUs and 96 vCPUs in total. Figure 6(d) shows that Ray Data is able to achieve near-perfect strong scaling.
Fault tolerance.
CPU executor failures are common in heterogeneous clusters mostly due to preprocessing tasks running out of memory. Ray Data can transparently recover from isolated executor failures without interrupting the job. To further evaluate the cost of fault recovery, we intentionally fail the CPU worker node after 10 min, and reconnecting it after another 1 min. We compare Ray Data’s native lineage reconstruction failure recovery against checkpointing, which is commonly used in stream processing systems. Figure 6(c) shows that Ray Data’s throughput drops to that of the remaining node during worker node failure, and restores when the node rejoins without interrupting the job. In comparison, we implement a checkpointing-based recovery, by saving progress every 5 min, then restarting the job to load from the last checkpoint when failure occurs. As expected, the system makes no progress until $t=18\mathrm{min}$, due to having to restart the job and perform redundant work.
Takeaways:
Ray Data enables scaling with heterogeneous clusters to reduce overall cost by 60%, despite using more nodes. Compared to stream processing systems that use global checkpointing, failures at the additional CPU-only nodes have little impact on end-to-end throughput, allowing further cost reduction via spot instances.
5.2 Training: Comparison to ML data loaders
5.2.1 ResNet Training
We run the ResNet-50 ImageNet training benchmark from MLPerf [26]. The data preprocessing pipeline loads images from local disk (local) or cloud storage (S3), decodes, and randomly crops and flips the images. We compare training throughput of tf.data vs. Ray Data on a g5.2xlarge VM. We do not measure PyTorch DataLoader, as tf.data showed comparable or better results for the same benchmark in [30].
Figure 8 shows training throughput over time. tf.data executes data preprocessing using a pool of worker threads running in each GPU trainer process. Thus, the job fate-shares with any preprocessing task that fails due to out-of-memory (OOM). When reading data from local disk, tf.data’s throughput is 19% lower than Ray Data’s because a lower batch size was required to prevent OOM failures. Meanwhile, Ray Data is able to complete because GPU trainer failures are isolated from CPU worker failures, and CPU workers can be respawned in seconds, without impacting pipeline throughput (§ 5.1.3).
When reading data from S3, tf.data is 88% slower than the max GPU throughput because S3 loading is the bottleneck. Meanwhile, Ray Data can use heterogeneous clusters to scale out S3 loading independent of the GPU trainers. By adding a m7i.2xlarge node for data preprocessing, the overall training throughput reaches 93% of the max GPU throughput.
Takeaways:
Compared to single-node ML data loaders, Ray Data offers: (1) failure isolation between heterogeneous resources, and (2) ability to leverage heterogeneous clusters.
5.2.2 Pre-Training Stable Diffusion
Pre-training of large ML models is one of the most demanding heterogeneous workloads. We run the Stable Diffusion (SD) pre-training pipeline shown in Figure 1(b) and compare different execution modes. We execute 1 training epoch over a dataset of 2 billion images, on a cluster of 4×4$\times$4 × p4de.4xlarge nodes, with 8 A100 GPUs on each node. This pipeline is challenging because it requires both CPUs and GPUs for data preprocessing: (1) loadText/loadImage+clip (CPU): Load pairs of image and text, perform preprocessing, (2) Encoder (GPU): Use a pre-trained encoder model, one for images and one for text, to produce dense embeddings, and (3) UNet.train() (GPU): Train SD on the embeddings.
Figure 8 shows training throughput and total cost. PyTorch DL is a data loader custom-built for PyTorch that statically partitions work on a process pool. Its throughput is lowest because Encoder preprocessing competes with trainers for GPU memory. Ray Data-staged emulates batch processing by running data preprocessing as an offline job and storing precomputed embeddings in cloud storage. This is preferable if the same embeddings are used multiple times. Ray Data-staged achieves 19% higher throughput because UNet is given full GPUs.
Ray Data runs all data preprocessing concurrently with training. This is preferable if using random transforms or for iterative development. Ray Data also leverages heterogeneous clusters, placing Encoders on smaller A10G GPUs (g5.2xlarge). This results in 31% better throughput than PyTorch DL, because UNet has full GPU resources, and 15% better throughput than Ray Data-staged, because embeddings are kept in memory.
Takeaways:
Compared to existing ML data loaders, Ray Data can: (1) be used for both batch and online data preprocessing, and (2) leverage heterogeneous distributed execution.
Figure 9: Synthetic benchmark run times for systems under different memory limits. Grey means the system is unable to finish due to OOM. Ray Data(-Part.) means Ray Data without streaming repartition. Ray Data(-Adapt.) means Ray Data without adaptive memory-aware scheduling.
Figure 10: Effect of partition sizes on throughput in Ray Data.
5.3 Microbenchmarks
5.3.1 Memory-aware scheduling
We evaluate how batch, stream, and streaming batch systems schedule heterogeneous pipelines with and without memory pressure. The 3-stage pipeline is: (1) Load (CPU): 160 tasks, each producing 500 1 MB rows after 5s, (2) Transform (CPU): sleep for 0.5s per row, then return a different 1 MB row, (3) Inference (GPU): 0.5s per batch of 100 rows. We use 1 m6i.2xlarge node with 8 vCPUs, 4 simulated GPU slots, and 32 GB RAM. The theoretical best job completion time with unlimited memory is $(160\times 5s+800\times 0.5s)/8=150s$.
Figure 10 shows job completion time vs. total memory limit. We limit memory through system-specific configurations, e.g., executor memory for Spark. We also use POSIX rlimit to verify that each system respects its memory limit, and tune each system’s parallelism (e.g., executor count) if not.
Spark materializes all data between stages, achieving at best 2.35×2.35$\times$2.35 × optimal run time. At 12–14GB memory, Spark must use fewer executors resulting in 4.34×4.34$\times$4.34 × the optimal run time, and at lower memory limits, Spark is unable to finish. This is because Spark requires static partitioning (§ 2.2), and the initial number of Load tasks produces too-large partitions.
Flink is less sensitive to the memory limit than Spark and achieves up to 1.68×1.68$\times$1.68 × optimal. This is because executors dynamically materialize output partitions to avoid running out of memory (§ 2.3). At lower memory limits, Flink must run fewer executors because it uses multithreading and slot sharing to multiplex a CPU slot among physical operators, making executors vulnerable to OOM under memory pressure. This results in up to 2×2$\times$2 × worse throughput.
We also compare against tf.data because unlike PyTorch DL, it offers an adaptive scheduler and memory budget, similar to Ray Data. However, we found that the memory budget was not always enforced, requiring manual tuning of the thread count. tf.data achieves the same throughput as Ray Data at 16 GB memory limit, but is unable to finish at lower memory limits.
Ray Data is able to finish in 1.3×1.3$\times$1.3 × the optimal run time at all memory limits except the lowest, due to Ray Data’s streaming repartition (Figure 3(b)) and adaptive scheduler (§ 4.3). We further conduct ablation studies on Ray Data’s performance. Ray Data(-Part.) disables Ray Data’s streaming repartition, resulting in too-large initial partitions similar to Spark. Ray Data(-Adapt.) disables Ray Data’s adaptive scheduler, resulting in a conservative policy similar to Figure 4(a) and 10–88% worse performance than Ray Data. Ray Data is also less sensitive than Flink to memory pressure because the system explicitly time-slices executors at task granularity, instead of using multithreading.
Takeaways:
For heterogeneous applications under memory pressure, batch processing systems are unstable. Ray Data is as stable as Flink, due to its streaming repartition, and also more adaptive, thanks to its scheduler.
5.3.2 Overhead of partitioning
Compared to stream processing, Ray Data also uses dynamically sized partitions, but with a centralized scheduler. A possible concern is the system overhead per partition. We evaluate the impact of partition number on throughput in Ray Data with a 2-stage synthetic pipeline. We use 8192×8192$\times$8192 ×1 MB input rows, and simulate 10 ms processing time per row per stage. Figure 10 shows the throughput vs. partition size. The smallest partition sizes incur overhead from RPCs and bookkeeping, and the largest result in poor load-balancing. To strike a balance, Ray Data’s default target partition size is 128 MB.
6 Related Work
Unifying batch and stream processing systems.
Recent efforts to unify the batch and stream processing models include Apache Beam [4] and Google Cloud Dataflow [3]. They focus on providing a unified API layer, rather than unifying the execution model. Attempts at execution model unification include Naiad [28], Flink’s batch execution mode [9], and Spark Streaming [47]. Naiad shows that the stream processing model is suitable for producing results both incrementally and in bulk, but it does not support dynamic parallelism reconfiguration. Like other batch processing systems, Flink’s batch execution mode executes one stage at a time.
Spark Streaming partitions the input stream into discretized “microbatches”, each executed as a distinct Spark job. Drizzle [43] improves on Spark Streaming’s latency without sacrificing adaptability, but suffers the same flaws for heterogeneous pipelines: pipelining across heterogeneous resources would require a custom inter-job Spark scheduler, and the data partitioning within a microbatch is static. Also, since Spark requires stateless tasks, it imposes high overheads in ML inference and training from repeatedly loading the model into GPU memory for each task. Petastorm [17] is a Python library that bridges Spark’s data processing capabilities to ML training frameworks. It shares the same limitations as Spark.
MillWheel [2] is a stream processing system that offers efficient reconfiguration and failover by combining decentralized physical logging with a centralized load-balancer off the critical path. It offers sophisticated APIs for real-time processing, including timers, watermarks, etc. In contrast, Ray Data targets offline processing, uses lineage-based recovery to avoid data logging, and the centralized scheduler dispatches all tasks for a global view and finer control over resources.
Other systems have explored building distributed data processing frameworks on top of task-parallel systems such as Ray [27, 44], CIEL [29] and Dask [37]. However, all of these systems are CPU-centric and do not consider memory or heterogeneous compute.
Scheduling for resource heterogeneity.
The scheduling problem described in Section 4.3 is most similar to the generalized processor sharing [33] problem. Our solution is inspired by the weighted fair queueing algorithm [6, 12]. The differences are (1) the flows in network scheduling are independent of each other, whereas operators in a data pipeline have dependencies, (2) multiple resource types, and (3) the packet processing time is usually fixed, whereas data operator processing times are unpredictable. (1) is important in that operators may produce significant intermediate data, which may not be released until the downstream operator executes.
Recent scheduling works attempt to adapt fair queuing and autotuning to heterogeneous resource environments. Dominant resource fair queuing [15] addresses the problem of (2) but not (1) or (3). tf.data [30] introduces an autotuning algorithm that uses gradient descent to find the parallelism for each operator that reduces end-to-end latency. We instead aim to maximize overall throughput; the partition size may be used to adjust end-to-end latency.
Streaming data loaders for ML.
Both PyTorch [34] and TensorFlow (tf.data [30]) provide data loaders optimized for map-style transforms for ML training. tf.data uses multithreading while PyTorch DataLoader uses multiprocessing. tf.data automatically and deterministically shards the dataset, while PyTorch DataLoader requires the user to shard the dataset themselves. However, both are single-node systems colocated with a GPU trainer and share similar limitations: they cannot execute multi-node, dataset sharding must be done before execution, and the training job fate-shares with the data loader.
Cloud-native streaming data loaders include MosaicML Streaming [21] and DeepLake [18]. The common innovations in these libraries are a specialized data format optimized for tensor storage and querying, and streaming data loading directly from cloud storage. They use Python multiprocessing for execution, which can be swapped out for Ray Data. This would decouple data loading from ML training/inference processes, thus allowing more flexible control of CPU parallelism and memory buffers and scale-out to heterogeneous nodes.
7 Discussion
One benefit of building Ray Data as a Ray library rather than a monolithic system such as Spark or Flink is that since the Ray core is in C++, Ray Data can support many frontend languages, including the lingua franca for ML data loading and transformations, Python. In contrast, systems built on non-native languages show high overheads when a Python frontend is used (e.g., Flink in § 5.1.1). Second, modifying the Ray Data scheduler is convenient, as it is written in the frontend and does not require re-compiling Ray [25].
Ray Data enables another key opportunity in future data processing systems: dynamic query planning. In this work, we present an online scheduler, but still make certain planning decisions statically, including the number of input partitions (§ 4.1), and the user specifies the initial cluster shape. We envision a fully autotuning and autoscaling system that can cohesively re-plan the application and resize the cluster.
Conclusion.
With the rise of large language models, even modalities such as text that are traditionally not compute-intensive to load and process may now require expensive deduplication [31], GPU-based embedding computation, and joins with image or video data (Figure 1(b)). In addition, as the parallelism strategies used in inference and training pipelines become more complex, future data processing systems must also support more flexible APIs for sharding and sharing data.
In general, we believe that ML systems will continue to grow in the complexity of their data processing needs, as evidenced by trends such as test-time training [39, 14], retrieval-augmented generation [24], and multimodal models [40, 32]. To keep up with this demand, we must build more flexible, heterogeneity-aware, and scalable data processing systems.
Appendix A Additional Microbenchmarks
A.1 Fractional Parallelism
In this microbenchmark, we demonstrate that the streaming batch model can maximize the resource utilization when fractional parallelism is required. Consider a two-stage data pipeline, in which the first stage takes 1 second on average, and the second stage takes 2 seconds. Ideally, the operator parallelisms should be set as $2:1$ to balance the throughput. In traditional stream processing systems such as Flink, this is unattainable on a 8-CPU machine, because it requires setting the operator parallelisms to be 2.67 and 5.33, respectively. Since these systems allocate executors to operators statically, they cannot support fractional parallelism. In contrast, the streaming batch execution model allows Ray Data to multiplex executors for both stages dynamically during run time. The Ray Data scheduler can dynamically start a task for either stage in order to balance the throughput, effectively achieving a parallelism ratio of $2:1$ over time. Figure 11 shows that when comparing to a static allocation of 4–4 executors for each stage, the dynamic allocation increases the utilization of the execution slots, manifested as fewer bubbles in the schedule, and 19% faster job completion time.
Figure 11: Dynamically allocated executor slots can achieve fractional parallelism with better resource utilization (fewer bubbles).
Appendix B Solver for Discrete-time Scheduling
To verify the efficacy of the online scheduling algorithm, we develop a discrete-time simulation environment, in which tasks have fixed execution times, and a discrete-time solver that can find the optimal schedule to run a data pipeline, subject to specified resource constraints.
The input to the solver is a data pipeline, the total data size, and the resource constraints. The data pipeline is described as a chain of operators. Each operator processes data in tasks. Each task has an input size and an output size measured in number of partitions, and we assume each task has a known duration. Each task also has a resource requirement, e.g. 1 CPU or 1 GPU.
The total data size is also measured in number of partitions. The resource constraints describe how many execution slots are available for each resource type (CPU or GPU), and also has a memory buffer limit, indicating how many intermediate partitions in total can be stored in the temporary memory buffer.
Finally, the solver has a length limit, measured in time ticks, for any solution returned. This is such that the solution space is bounded.
B.1 Algorithm
The solution space is defined by the set of all possible execution states. The execution state consists of:
  • Time since the start.
  • The state of each executor, i.e. which operator task is running.
  • The state of the shared memory buffer, which is the number of partitions stored in the buffer.
  • The state of each operator, which is the number of pending tasks.
The solver starts from the initial state (time 0, all executors idle, buffer empty, and all tasks pending). For each state, it generates the next state by emulating the execution: advancing the tick, updating executor states, updating the progress of running tasks, updating the memory buffer, etc. The number of next states is determined by the size of the set of all possible scheduling actions, which is the power set of all possible scheduling primitives. A scheduling primitive would be “schedule the next task operator $i$ onto executor $j$.”
The solver runs a variation of the A* search algorithm to try to arrive at the first completion state (in which no more tasks are pending). In the priority queue, the states are sorted by the number of completed tasks, i.e. it prioritizes states that make further progress. The solver returns the optimal job completion time after all possible states are visited.
The naive search algorithm is not practical due to its high time complexity ($O((E\cdot T)^{N})$), where $N$ is the total number of tasks, $E$ the total number of executors, and $T$ the time limit). We use the following optimizations to bring the complexity down to $O(2^{N}\cdot T)$, making it more practical for large scheduling problems.
  • Symmetry of tasks and executors. We assign a canonical ordering of the executors, i.e. the first task always starts on the lowest-numbered executor. This gets rid of a large class of duplicate states, where the task timings are the same, except that they run on different executors.
  • Temporal equivalence. We notice that the optimal job completion time given an execution state at time $t$ is the same, regardless of its execution history before $t$. This means all states that arrive at the same task progress at time $t$ are equivalent. This is crucial for reducing the number of duplicate states, and in many cases, reduces the problem to polynomial time.
For the scheduling microbenchmark in Section 5.3.1, the solver finds the optimal schedule with a total run time of 153 seconds.